OpenStudy (anonymous):

In a box containing marbles, 78 are blue, 24 are orange, and the rest are green. If the probability of selecting a green marble is 2/5, how many green marbles are in the box?

Vocaloid (vocaloid):

hint: set up a proportion let x = number of green marbles green marbles/total marbles = 2/5 know where to go from there?

OpenStudy (anonymous):

ok but i don't have the total marbles

OpenStudy (anonymous):

\[\frac{g}{g+78+24}=\frac{2}{5} \]Solve the above for g, the number of green marbles. Mathematica came up with g = 68 .

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

Your welcome.